# Quantifying sunspot group nesting with density-based unsupervised clustering

### v1

@@ Abstract
Sunspot groups often emerge in spatial–temporal clusters, known as nests or complexes of activity. Quantifying how frequently such nesting occurs is important for understanding the organisation and recurrence of solar magnetic fields. We introduce an automated approach to identify nests in the longitude–time domain and to measure the fraction of sunspot groups that belong to them. The method combines a smooth representation of emergence patterns with a density-based clustering procedure, validated using synthetic solar-like cycles and corrected for variations in data density.
We apply this method to 151 years of sunspot-group observations from the Royal Greenwich Observatory Photoheliographic Results (RGO, 1874–1976) and Kislovodsk Mountain Astronomical Station (KMAS, 1955–2025) catalogues.Across all cycles and latitude bands, the mean nesting degree is $\langle D\rangle=0.61\pm 0.12$, implying that about 60 percent all sunspot groups emerge within nests. Nesting is strongest at mid-latitudes (10–20), and results from the two independent datasets agree in the period of overlap. The identified nests range from compact clusters to long-lived, drifting structures, offering new quantitative constraints on the persistence and organisation of solar magnetic activity.
@@ 1 Introduction
Sunspots appear in the solar photosphere as visible signatures of magnetic flux bundles emerging from the solar interior. These bundles typically form structures with positive and negative polarities, which outline the footpoints of rising magnetic flux loops (Lidia15; Weber23). The emergence of sunspot groups (SGs) is one of the primary indicators of solar magnetic activity at photospheric levels.
@@
It has long been recognised that sunspot groups do not emerge randomly in longitude. Instead, they tend to cluster in space and time, forming so-called nests or complexes of activity. Early statistical work by bec55 and later analyses of synoptic magnetograms by gai83 highlighted the recurrent nature of these emergence patterns (bec55 called them Fleckenherde, or ‘spot herds’). These clusters, which may persist for several solar rotations, were found to exhibit characteristic longitudinal motions that deviate from the local (differential) rotation profile in both directions and occasionally show diverging and converging behaviour over time.
@@
A systematic and quantitative treatment of nesting was initiated by cas86, who applied cluster analysis to the Greenwich sunspot group data. Their analysis revealed statistically significant nests typically spanning less than $30^{\circ}$ in longitude and persisting up to 15 solar rotations. bro90 extended this work using three-dimensional (time, longitude, and latitude) clustering analysis of Greenwich data, showing that at least one-third of all sunspot appearances belong to intrinsically physical clusters, with characteristic longitudinal and latitudinal spreads of $\sim 2^{\circ}$ and $\sim 1^{\circ}$, respectively, and lifetimes of 1–7 months. The patterns showed hierarchical organisation that were called ‘nested nests’ with component separations below $25^{\circ}$ longitude.
@@
SG nesting was also sometimes connected to the so-called active longitudes, which exhibit persistent activity patterns across multiple rotations with $180^{\circ}$ separation between hemispheric activity herds and flip-flop cycles of 3.7 years (Berdyugina03; Usoskin05; Usoskin07).
@@
Despite these efforts, the quantitative characterisation of nesting remains challenging (Pojoga02). Earlier analyses often relied on subjective criteria or methods that assumed specific cluster shapes or rotation profiles. Moreover, the increasing availability of long-term, homogeneous sunspot catalogues motivates the development of automated and reproducible techniques to quantify clustering behaviour. Modern data-driven machine learning techniques, including non-parametric density estimation and unsupervised pattern-recognition methods, offer new opportunities to revisit these questions (AsensioRamos23).
@@ 2.1 Data
We analyse sunspot group (SG) data from two complementary observational programs. Our primary dataset comprises the Royal Greenwich Observatory Photoheliographic Results (RGO, 1874–1976), providing systematic observations spanning Cycles 12–19 (Willis13). To extend coverage through the modern era, we use data from the Kislovodsk Mountain Astronomical Station (KMAS, 1955–2025), covering Cycles 19–25 (Nagovitsyn07).
@@
The datasets overlap during 1954–1976 (Cycles 19–20), enabling direct comparison and cross-calibration. Cross-calibration studies show that KMAS individual group areas are systematically smaller than RGO measurements by approximately 3%, with a scaling factor $b_{\rm RGO-KMAS}=1.031\pm 0.056$ (Mandal20). For our analysis, we apply this correction to bring KMAS areas to the RGO scale, ensuring homogeneity across the combined dataset. The overlap period also enables validation of our nesting methodology across independent observational programs (Section 3.2).
@@ 2.2 Methods
To evaluate SG-nesting quantitatively, we employ Kernel Density Estimation (KDE) for visualisation of emergence density patterns, combined with the DBSCAN (Density Based Spatial Clustering of Applications with Noise) clustering algorithm for discrete nest identification. The KDE provides intuitive visual guidance, while DBSCAN assigns each sunspot group to either an identified nest or classifies it as an isolated emergence. As the main quantity that quantifies nesting, we define the degree of nesting as the fraction of SGs in nests, derived from each latitude slice’s time window, which is a full solar cycle. It is expressed as
@@ 2.2.1 Kernel Density Estimation
Kernel Density Estimation (KDE) is a non-parametric technique for estimating the probability density function of a dataset by placing a selected kernel function $K_{H}$ at each data point and summing their contributions. For a dataset of $n$ SGs with emergence longitudes and times $(\phi_{i},t_{i})$ and maximum observed areas $A_{i}$, we define area-weighted weights as $w_{A,i}=\ln(A_{i})/\ln(A_{\rm max})$, where $A_{\rm max}=\max_{j}(A_{j})$ is the largest SG area in the analysis window. The two-dimensional area-weighted kernel density estimator in the longitude-time space is
@@
where $K_{H}$ is a bivariate Gaussian kernel with bandwidth parameters $h_{\phi}$ and $h_{t}$ determined using Scott’s rule (Scott79). Prior to KDE, both longitude and time coordinates are standardised using z-score normalisation. The area weighting ensures that larger SGs contribute more significantly to the density estimation, reflecting their enhanced magnetic flux content.
@@ 2.2.2 Clustering with DBSCAN
To identify discrete nesting patterns in SG emergence, we employ the Density Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester96). Unlike traditional clustering methods such as k-means, DBSCAN does not require specifying the number of clusters beforehand and can identify clusters of arbitrary shapes, making it well-suited for detecting the elongated or irregular nesting patterns expected in solar active region emergence.
@@ 2.2.3 Parameter Selection and Validation
$\varepsilon_{\rm eff}=\varepsilon_{0}\times\min\left(1.5,\sqrt{\frac{N_{\rm ref}}{N}}\right),$ (3)
where $\varepsilon_{0}=0.11$ is the baseline parameter, $N$ is the number of SGs in the analysis window, and $N_{\rm ref}=150$ is a reference sample size. The reference value was determined empirically: synthetic tests showed that windows with $N\gtrsim 150$ yield stable recovery of nests without excessive fragmentation, so the correction compensates only in sparser cases. This correction increases $\varepsilon$ by up to 50% for sparse datasets, preventing excessive fragmentation of genuine nesting structures during low-activity periods. For dense data slices with $N>N_{\rm ref}$, no correction is applied ($\varepsilon_{\rm eff}=\varepsilon_{0}$).
@@ 2.2.5 Statistical Significance
$p(X\geq n_{c}|N_{\rm expected})=1-F_{\rm Poisson}(n_{c}-1;\lambda),$ (5)
@@ 3.2 Cross-Validation Between Datasets
To assess the robustness of our methodology and quantify systematic differences between observational programs, we applied our nesting analysis independently to both RGO and KMAS data during their overlap period (1954–1976, covering Cycles 19–20). Both datasets were processed using identical analysis windows and latitude bands ($5^{\circ}$ intervals), with KMAS areas scaled to the RGO reference frame using $b_{\rm RGO-KMAS}=1.031$ (Mandal20).
@@
with RMS scatter of 0.158. The slope less than unity indicates that the offset between datasets increases with nesting degree: at low $D$ values, the difference is smaller ($\sim 0.05$ at $D_{\rm KMAS}=0.1$), while at high $D$ values it becomes substantial ($\sim 0.15$ at $D_{\rm KMAS}=0.7$). This pattern is consistent across both cycles (Cycle 19: $r=0.891$; Cycle 20: $r=0.657$) and both hemispheres (N: mean bias +0.140+0.140; S: mean bias +0.106+0.106).
RGO’s superior completeness for smaller SGs enables better detection of weak or dispersed nesting structures, particularly important when overall nesting degree is high and many small groups cluster around large active regions. Figure 4 illustrates this effect for a representative window (Cycle 19, $15^{\circ}$$20^{\circ}$ latitude): RGO identifies substantially more groups, resulting in higher recovered nesting degrees in both hemispheres ($D_{\rm RGO}=0.768$ vs. $D_{\rm KMAS}=0.663$ for the northern band).
@@ 3.3 Spatio-temporal Organisation of Nesting
12 1878–1890 RGO 6 42.2 0.0 0.552 0.128
13 1890–1902 RGO 7 54.6 0.0 0.539 0.151
14 1902–1913 RGO 6 52.5 0.0 0.560 0.095
15 1913–1923 RGO 7 64.4 0.0 0.593 0.178
16 1923–1933 RGO 7 62.6 0.0 0.610 0.081
17 1933–1944 RGO 7 68.3 0.0 0.587 0.158
18 1944–1954 RGO 7 68.4 0.0 0.590 0.181
19 1954–1964 RGO 7 79.9 0.0 0.643 0.092
KMAS 7 64.9 0.0 0.684 0.070
20 1964–1976 RGO 7 70.7 0.0 0.640 0.072
KMAS 7 60.6 0.0 0.661 0.087
21 1976–1986 KMAS 7 61.4 0.0 0.671 0.103
22 1986–1996 KMAS 7 52.7 0.0 0.642 0.072
23 1996–2008 KMAS 7 51.3 0.0 0.652 0.091
24 2008–2019 KMAS 6 49.3 0.0 0.664 0.063
25 2019–2025 KMAS 7 45.3 0.0 0.627 0.101
@@ 3.3.2 Longitude Distribution
This longitude independency distinguishes our full-cycle analysis from shorter timescale studies of active longitudes. On rotational timescales ($\lesssim 1$ year), pronounced longitude preferences are commonly observed (Berdyugina03; Usoskin07). However, these structures evidently lack the multi-year persistence necessary to generate cycle-averaged longitude asymmetries in nesting patterns.
@@ 3.4 Correlation with Solar Activity
The trend is not strictly linear and there is considerable cycle-to-cycle scatter, especially in the KMAS dataset where small-number statistics affect low- and high-activity edges. Nevertheless, the overall behaviour suggests that more coherent toroidal flux systems during strong cycles give rise to more frequent or more pronounced longitudinal clustering This finding is consistent with the enhanced large-scale magnetic organisation expected during periods of elevated solar activity.

### v2

@@ Abstract
Sunspot groups often emerge in spatial–temporal clusters, known as nests or complexes of activity. Quantifying how frequently such nesting occurs is important for understanding the organisation and recurrence of solar magnetic fields. We introduce an automated approach based on kernel density estimation and DBSCAN clustering to identify nests in the longitude–time domain and to measure the fraction of sunspot groups that belong to them. The method combines a smooth representation of emergence patterns with a density-based clustering procedure, validated using synthetic solar-like cycles and corrected for variations in data density.
We apply this method to 151 years of sunspot-group observations from the Royal Greenwich Observatory Photoheliographic Results (RGO, 1874–1976) and Kislovodsk Mountain Astronomical Station (KMAS, 1955–2025) catalogues.Across all cycles and latitude bands, the mean nesting degree is $\langle D\rangle=0.61\pm 0.12$, implying that about 60 percent all sunspot groups emerge within nests. Nesting is strongest at mid-latitudes (10–20), and results from the two independent datasets agree in the period of overlap. The nesting degree significantly correlates with the solar activity level, with the correlation strengthening when small groups are excluded. The characteristic inter-nest spacing contracts from $\sim$200-500 Mm at low activity to $\sim$60-100 Mm at solar maximum, approaching typical sunspot-group dimensions. The identified nests range from compact clusters to long-lived, drifting structures, offering new quantitative constraints on the persistence and organisation of solar magnetic activity.
@@ 1 Introduction
Sunspots appear in the solar photosphere as visible signatures of magnetic flux bundles emerging from the solar interior. These bundles typically form structures with positive and negative polarities, which outline the footpoints of rising magnetic flux loops (van Driel-Gesztelyi and Green, 2015; Weber et al., 2023). The emergence of sunspot groups (SGs) is one of the primary indicators of solar magnetic activity at photospheric levels.
@@
It has long been recognised that sunspot groups do not emerge randomly in longitude. Instead, they tend to cluster in space and time, forming so-called nests or complexes of activity. Early statistical work by Becker (1955) and later analyses of synoptic magnetograms by Gaizauskas et al. (1983) highlighted the recurrent nature of these emergence patterns (Becker 1955 called them Fleckenherde, or ‘spot herds’). These clusters, which may persist for several solar rotations, were found to exhibit characteristic longitudinal motions that deviate from the local (differential) rotation profile in both directions and occasionally show diverging and converging behaviour over time.
@@
A systematic and quantitative treatment of nesting was initiated by Castenmiller et al. (1986), who applied cluster analysis to the Greenwich sunspot group data. Their analysis revealed statistically significant nests typically spanning less than $30^{\circ}$ in longitude and persisting up to 15 solar rotations. Brouwer and Zwaan (1990) extended this work using three-dimensional (time, longitude, and latitude) clustering analysis of Greenwich data, showing that at least one-third of all sunspot appearances belong to intrinsically physical clusters, with characteristic longitudinal and latitudinal spreads of $\sim 2^{\circ}$ and $\sim 1^{\circ}$, respectively, and lifetimes of 1–7 months. The patterns showed hierarchical organisation that were called ‘nested nests’ with component separations below $25^{\circ}$ longitude.
Benevolenskaya (2005) examined activity nests within the broader context of complexes of solar activity, linking their formation and evolution to the topology of the non-axisymmetrical component of the solar magnetic field. Her analysis connected nesting patterns to large-scale coronal structures and suggested that activity complexes may have roots deep in the tachocline (see also Raphaldini et al., 2023).
@@
SG nesting was also sometimes connected to the so-called active longitudes, which exhibit persistent activity patterns across multiple rotations with $180^{\circ}$ separation between hemispheric activity herds and flip-flop cycles of 3.7 years (Berdyugina and Usoskin, 2003; Usoskin et al., 2005, 2007).
@@
Despite these efforts, the quantitative characterisation of nesting remains challenging (Pojoga and Cudnik, 2002). Earlier analyses often relied on subjective criteria or methods that assumed specific cluster shapes or rotation profiles. Moreover, the increasing availability of long-term, homogeneous sunspot catalogues motivates the development of automated and reproducible techniques to quantify clustering behaviour. Modern data-driven machine learning techniques, including non-parametric density estimation and unsupervised pattern-recognition methods, offer new opportunities to revisit these questions (Asensio Ramos et al., 2023).
@@ 2.1 Data
We analyse sunspot group (SG) data from two complementary observational programs. Our primary dataset comprises the Royal Greenwich Observatory Photoheliographic Results (RGO, 1874–1976), providing systematic observations spanning Cycles 12–19 (Willis et al., 2013). To extend coverage through the modern era, we use data from the Kislovodsk Mountain Astronomical Station (KMAS, 1955–2025), covering Cycles 19–25 (Nagovitsyn et al., 2007).
@@
The datasets overlap during 1954–1976 (Cycles 19–20), enabling direct comparison and cross-calibration. Cross-calibration studies show that KMAS individual group areas are systematically smaller than RGO measurements by approximately 3%, with a scaling factor $b_{\rm RGO-KMAS}=1.031\pm 0.056$ (Mandal et al., 2020). For our analysis, we apply this correction to bring KMAS areas to the RGO scale, ensuring homogeneity across the combined dataset. The overlap period also enables validation of our nesting methodology across independent observational programs (Section 3.2).
@@ 2.2 Methods
To evaluate SG-nesting quantitatively, we employ Kernel Density Estimation (KDE) for visualisation of emergence density patterns (Scott, 1979), combined with the DBSCAN (Density Based Spatial Clustering of Applications with Noise) clustering algorithm (Ester et al., 1996) for discrete nest identification. The KDE provides intuitive visual guidance, while DBSCAN assigns each sunspot group to either an identified nest or classifies it as an isolated emergence. As the main quantity that quantifies nesting, we define the degree of nesting as the fraction of SGs in nests, derived from each latitude slice’s time window, which is a full solar cycle. It is expressed as
@@ 2.2.1 Kernel Density Estimation
Kernel Density Estimation (KDE) is a non-parametric technique for estimating the probability density function of a dataset by placing a selected kernel function $K_{H}$ at each data point and summing their contributions (Scott, 1979). For a dataset of $n$ SGs with emergence longitudes and times $(\phi_{i},t_{i})$ and maximum observed areas $A_{i}$, we define area-weighted weights as $w_{A,i}=\ln(A_{i})/\ln(A_{\rm max})$, where $A_{\rm max}=\max_{j}(A_{j})$ is the largest SG area in the analysis window. The two-dimensional area-weighted kernel density estimator in the longitude-time space is
@@
where $K_{H}$ is a bivariate Gaussian kernel with bandwidth parameters $h_{\phi}$ and $h_{t}$ determined using Scott’s rule (Scott, 1979). Prior to KDE, both longitude and time coordinates are standardised using z-score normalisation. The area weighting ensures that larger SGs contribute more significantly to the density estimation, reflecting their enhanced magnetic flux content.
@@ 2.2.2 Clustering with DBSCAN
To identify discrete nesting patterns in SG emergence, we employ the Density Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester et al., 1996). Unlike traditional clustering methods such as k-means, DBSCAN does not require specifying the number of clusters beforehand and can identify clusters of arbitrary shapes, making it well-suited for detecting the elongated or irregular nesting patterns expected in solar active region emergence.
@@ 2.2.3 Parameter Selection and Validation
$\varepsilon_{\mathrm{eff}}=\left\{\begin{array}[]{ll}\varepsilon_{0}\times\min\left(1.5,\sqrt{N_{\mathrm{ref}}/N}\right),&N<N_{\mathrm{ref}}\\ \varepsilon_{0},&N\geq N_{\mathrm{ref}}\end{array}\right$. (5)
where $\varepsilon_{0}=0.11$ is the baseline parameter, $N$ is the number of SGs in the analysis window, and $N_{\rm ref}=350$ is a reference sample size. The reference value was determined empirically: synthetic tests showed that windows with $N\gtrsim 350$ yield stable recovery of nests without excessive fragmentation, so the correction compensates only in sparser cases. This correction increases $\varepsilon$ by up to 50% for sparse datasets, preventing excessive fragmentation of genuine nesting structures during low-activity periods. For dense data slices with $N>N_{\rm ref}$, no correction is applied ($\varepsilon_{\rm eff}=\varepsilon_{0}$).
@@ 2.2.5 Statistical Significance
$p(X\geq n_{c}|N_{\rm expected})=1-F_{\rm Poisson}(n_{c}-1;N_{\rm expected}),$ (7)
@@ 3.2 Cross-Validation Between Datasets
To assess the robustness of our methodology and quantify systematic differences between observational programmes, we applied our nesting analysis independently to both RGO and KMAS data during their overlap period (1954–1976, covering Cycles 19–20). Both datasets were processed using identical analysis windows and latitude bands ($5^{\circ}$ intervals), with KMAS areas scaled to the RGO reference frame using $b_{\rm RGO-KMAS}=1.031$ (Mandal et al., 2020).
@@
with RMS scatter of 0.158. The positive intercept implies that RGO systematically yields higher nesting degrees than KMAS over most of the observed range. The slope below unity indicates that the magnitude of this offset decreases with increasing nesting degree ($\sim 0.28$ at $D_{\rm KMAS}=0.1$ and $\sim 0.07$ at $D_{\rm KMAS}=0.7$). This pattern is consistent for both cycles (Cycle 19: $r=0.891$; Cycle 20: $r=0.657$) and both hemispheres (N: mean bias +0.140+0.140; S: mean bias +0.106+0.106).
The weaker correlation in Cycle 20 (r = 0.657) compared to Cycle 19 (r = 0.891) suggests that systematic differences between the datasets might depend on the average activity level. During lower-activity periods (Cycle 20 as compared to 19), RGO’s superior small-group detection may become more important for identifying weak or dispersed nests. Alternatively, this difference may partly reflect changes in KMAS observing procedures during its early operational phase following its start in 1955.
This completeness effect remains important even at high nesting degrees, where numerous small groups cluster around large active regions. Figure 4 illustrates this effect for a representative window (Cycle 19, $15^{\circ}$$20^{\circ}$ latitude): RGO identifies substantially more groups, resulting in higher recovered nesting degrees in both hemispheres ($D_{\rm RGO}=0.768$ vs. $D_{\rm KMAS}=0.663$ for the northern band).
@@ 3.3 Spatio-temporal Organisation of Nesting
12 1878–1890 RGO 6 42.2 24.6 0.552 0.128
13 1890–1902 RGO 7 54.6 39.7 0.539 0.151
14 1902–1913 RGO 6 52.5 34.8 0.560 0.095
15 1913–1923 RGO 7 64.4 45.1 0.593 0.178
16 1923–1933 RGO 7 62.6 43.5 0.610 0.081
17 1933–1944 RGO 7 68.3 44.3 0.587 0.158
18 1944–1954 RGO 7 68.4 46.3 0.590 0.181
19 1954–1964 RGO 7 79.9 37.3 0.643 0.092
KMAS 7 64.9 37.6 0.684 0.070
20 1964–1976 RGO 7 70.7 33.1 0.640 0.072
KMAS 7 60.6 37.3 0.661 0.087
21 1976–1986 KMAS 7 61.4 39.4 0.671 0.103
22 1986–1996 KMAS 7 52.7 33.5 0.642 0.072
23 1996–2008 KMAS 7 51.3 27.7 0.652 0.091
24 2008–2019 KMAS 6 49.3 30.0 0.664 0.063
25 2019–2025 KMAS 7 45.3 33.3 0.627 0.101
@@ 3.3.2 Longitude Distribution
This longitude independency distinguishes our full-cycle analysis from shorter timescale studies of active longitudes. On rotational timescales ($\lesssim 1$ year), pronounced longitude preferences are commonly observed (Berdyugina and Usoskin, 2003; Usoskin et al., 2007; Korsós et al., 2024). However, these structures evidently lack the multi-year persistence necessary to generate cycle-averaged longitude asymmetries in nesting patterns.
@@ 3.4 Correlation with Solar Activity
The trend is not strictly linear and there is considerable cycle-to-cycle scatter, especially in the KMAS dataset where small-number statistics affect low- and high-activity edges. Nevertheless, the overall behaviour suggests that more coherent toroidal flux systems during strong cycles give rise to more frequent or more pronounced longitudinal clustering. This finding is consistent with the enhanced large-scale magnetic organisation expected during periods of elevated solar activity.
@@
For each threshold, we recomputed nesting degrees for all time-latitude windows and recalculated the correlation with activity level.
3.4.1 Correlation stability
One key question is whether the correlation between nesting degree and activity level persists when the analysis is restricted to larger groups, or whether it arises primarily because high-activity periods include more groups available to form apparent clusters. To test this, we repeated the analysis for progressively increasing minimum area thresholds.
Figure 9 (left panel) shows the correlation coefficient between nesting degree and the activity level (total area across time-latitude windows) as a function of the area threshold. For thresholds up to 200 MSH (KMAS) and 300 MSH (RGO), both datasets maintain strong, highly significant correlations (KMAS: r=0.70r=0.70-$0.79$; RGO: r=0.60r=0.60-$0.78$, all $p<10^{-12}$). Notably, in RGO the correlation increases from $r=0.60$ at zero threshold to $r=0.78$ at about 200 MSH, indicating that the nesting–activity relationship is strongest for medium-to-large groups. This behaviour argues against a statistical counting artefact: if the correlation were driven mainly by abundant small groups during high-activity periods, excluding them would weaken the correlation. Instead, the correlation strengthens, suggesting that the organisation of medium-to-large active regions ($\sim$100-300 MSH) plays a dominant role. At thresholds above 300 MSH, the correlation decreases as sample sizes become small.
Figure 9: Pearson correlation coefficients between the nesting degree and the activity level (left) and the nesting degree (right) as a function of the minimum area threshold, for KMAS (blue) and RGO (red) datasets. Star symbols denote correlations with $p<0.01$. The green region shows the region where the nesting-activity correlation are rather stable. The error bars show $1\sigma$ intervals.
3.4.2 Nesting Hierarchy
Beyond validating the nesting-activity correlation, the threshold analysis in Sect. 3.4.1 reveals hierarchical organisation in SG emergence. Figure 9 (right panel) shows mean nesting degree as a function of the area threshold for both datasets. Both KMAS (1955–2025, Cycles 19–25) and RGO (1874–1976, Cycles 11–20) exhibit consistent patterns: mean nesting degree decreases from $\sim$0.50–0.60 for all groups to $\sim$0.23–0.27 for groups exceeding 400–450 MSH. This decline indicates that small SGs are preferentially associated with clustered emergence, whereas large regions alone exhibit weaker clustering.
The non-zero asymptotic nesting degree of $\sim$0.25 at high thresholds indicates that even large active regions retain some tendency to form nests. The substantially higher nesting degree at low thresholds ($\sim$0.50) reflects ‘parasitic’ flux emergence, where small groups preferentially emerge near existing large active regions, possibly led by fragmentation of large-scale toroidal flux loops into smaller parts, clustering around the nest centre.
Overall, these results are consistent with a hierarchical emergence pattern: (1) small, parasitic groups (<<200 MSH) preferentially emerge near large active regions, producing strong clustering, and (2) large active regions themselves (>>300 MSH) emerge with reduced but persistent longitudinal organisation (see also Jiang et al., 2011, Sect. 3.3).
[…]
3.5 Inter-nest Length Scale
To characterise the spatial organisation of nesting, we compute an inter-nest length scale for each analysis window ($5^{\circ}$ latitude band over a full solar cycle),
$L_{\rm typ}=(A/N_{\rm sig})^{1/2},$ (9)
where $A=360^{\circ}\times 5^{\circ}$ is the total surface area of the latitude band, and $N_{\rm sig}$ is the number of statistically significant nests identified in that window. This metric represents the mean separation obtained by uniformly distributing $N_{\rm sig}$ nests across the available longitude range. We convert this angular separation to physical distance (Mm) using a reference latitude of $15^{\circ}$ ($\cos(15^{\circ})\simeq 0.966$), representative of the mid-latitude active region belt where most nesting occurs.
Figure 10: Inter-nest length scale $L_{\rm typ}$ as a function of the overall activity level in terms of the total SG area in each window considered. The range of SG longitudinal (E-W) separations is marked with a green band, where the gradient qualitatively indicates the size-dependent emergence likelihood.
Figure 10 shows the inter-nest length scale as a function of total SG area (a proxy for solar activity level) for both RGO (1874–1976) and KMAS (1954–2025) latitude-time windows. Both datasets exhibit a strong negative correlation between $L_{\rm typ}$ and activity level, with Pearson coefficients 0.647-0.647 ($p\sim 10^{-15}$) for RGO and 0.682-0.682 ($p\sim 10^{-13}$) for KMAS. At low activity levels (sparse emergence), $L_{\rm typ}$ reaches 200-500 Mm, while at high activity levels it converges to $\sim$60-100 Mm, comparable to typical SG elongations in the E-W direction (3–10, corresponding to $\sim$35–117 Mm at $15^{\circ}$ latitude). The mean value of $L_{\rm typ}$ is approximately 125 Mm for both datasets.
However, $L_{\rm typ}$ represents a phase–space density measure that integrates over the full solar cycle duration, including periods of minimal or no emergence. Since nests preferentially cluster in active longitude zones during emergence (e.g., Gaizauskas et al., 1983), the actual local spacing between nests within these zones is likely substantially smaller than $L_{\rm typ}$. Nevertheless, the strong anti-correlation with activity level demonstrates that nests pack more compactly in phase space during solar maximum, with the convergence to SG elongation scales suggesting dense spatial organisation during peak activity.